# Improving Technical ”How-to” Query Accuracy with Automated Search Results Verification and Reranking

### v1

@@ 3.2.1. Runtime UI Context Detection
This module is designed to gather data about the device’s current User Interface (UI) context. The UI context includes a screenshot of the UI and metadata of ”actionable regions”. An ”actionable region” refers to a region on the screen where actions, such as clicking or scrolling, can be performed. These regions are hierarchically organized and can be associated with a textual description. For example, a region could be a navigation bar, with a back button and title as embedded child regions.
Sometimes, a region doesn’t contain any textual description, but child regions do have more meaningful textural information, which falls into the missing information in terms of action prediction. We use a control distillation process to fill in this missing information, depicted in Figure 3. First, invisible or non-actionable regions are filtered out. Then, descriptions from the nearest non-actionable child regions are consolidated to fill the missing information.
3.2.2. Multi-modal Action Prediction Model
Following Yang et al. (2023b); Zhang and Zhang (2023); Sun et al. (2022), the Action Prediction Model prompts GPT4-V OpenAI et al. (2023); Yang et al. (2023a), where the current UI context and multi-modal action chains (history of UI screenshots paired with each action taken), are passed as input to yield target action from a set of candidate actions333Injecting the UI context into the prompt can help model determine the next action, but if we can use it as an input to train or finetune the model, we expect more accurate action prediction results..
For timestep t , we try to predict the next action to take by prompting GPT4-V. A prompt is constructed using the following parameters and passed into GPT4-V:
@@ 3.2.2. Multi-modal Action Prediction Model
: List of screenshots of UI screens from previous actions taken.
: Action description for each previous action taken. Each previous action includes both the action name and the chosen actionable region description.
: Set of possible actions based on all kinds of actions currently found in .
: User attributes that aren’t covered in instructions, such as a sample username, password, and location.
@@ 3.2.3. Execution Proxy
It takes a generated action and its parameters returned from the Action Prediction Model, locates the action to the corresponding region, finally executes the action.
@@
It facilitates the execution information of agents to other modules, helping us analyze and monitor the execution progress.
@@ 3.3. Reranking
Execution information that include both textual and visual information (UI screenshot and corresponding actionable regions) is collected as listed in table 1. A feature extraction module converts these raw data into a reranking feature vector. To avoid unnecessary complexity initially, we now use statistical measures of execution information, and left as future work to explore more advanced feature generation techniques, such as time series transformer or feature learning.
@@ 3.3.1. Feature Design
Commonly, all textual and visual features are normalized to a float value between 0 and 1. Below is a brief description of our used features.
@@
Q is the set of tokens from the search query removed of stopwords, where is the term frequency of token in the extracted instructions I.
@@
: Average term frequency of common keywords in instructions: This feature measures the average frequency of common keywords found in the extracted instructions. We have extracted instructions from about 1000 web pages from Google Android Help Website and identified a set of most common keywords, removing stopwords.
@@
: ”Instructions completion degree” measures the degree to which the extracted instructions have been completed. GPT4-V is prompted to return a completion score ranging from [0, 1], passing the history of UI screenshots for each action taken during execution and extracted instructions as input.
: average, min, max, variance of Action term ratio in instructions measures the average, minimal, maximum, standard deviation of the percentage of instruction words that appear in each action description during the entire execution.
: average, min, max, variance of Visible UI Term Ratio in instructions measures the average, minimal, maximum, standard deviation of the percentage of the current screen UI control’s visible texts appeared in the instructions during the entire execution. This metric indicates the alignment between the current UI screen and the instructions. For example, at the sign-up phase, instructions often reference fields like username, password, and gender. A higher correspondence of these terms with UI elements suggests that the agent is progressing correctly.
: average, min, max, variance of Distilled UI ratio in instructions describes the average, minimal, maximum, standard deviation of the occurrence of distilled UI control text in the instructions during the entire execution. These metrics reflect the likelihood of our execution model accurately selecting the appropriate UI control from all distilled options.
to : These capture how much of each extracted step is executed by comparing the list of actions taken for a given step444A given step from a plan can require multiple actions to complete to its step text.
@@
Due to the page limit, we leave the exact equations for calculating the heuristic features to in our open source web site (TBA).
@@ 3.3.2. Reranking Models
In this section, we will illustrate which candidate pages will be extracted out reranking features and the details of our reranking models.
@@
When the system fails to extract instructions or those pages do not contain any instruction, a candidate page does not progress to the on-device execution stage. Therefore, our reranking module only uses a heuristic to prioritize all verified pages and rerank them subsequently. In our experiments, a page is considered as verified only if it progresses to stage 2: that is, the execution module can successfully execute at least one step and the predicted instruction completion rate is non-zero. All other Pages are kept with their original relative order (i.e. not reranked).
@@
Without losing the generality, three different learning-to-rank(LTR) models can be trained to rerank the verified pages using pre-defined features.
The first model is a simple pairwise LTR logistic regression model, which optimizes the relative order of pairs of pages. Another two types of models are multilayer perception (MLP) transformer models, short for TMLP, driven by two loss functions. One is LambdaLoss (Wang et al., 2018), a pairwise loss function. For each pair of items, the lambda value represents the change in the overall ranking metric, if the order of the two items were swapped. The lambda values are used to weight the errors in the pairwise comparisons, thus this loss function will penalize misorderings more if they have a higher lambda value. Another is NeuralNDCG loss (Qin et al., 2020), a method directly optimizing NDCG metric. This method introduces a NeuralSort operator based on a differentiable function (softmax) to approximate the non-differentiable ground-truth sort operator.
@@ 4. Experiments
Experiments are set up to analyze whether the proposed method can improve the retrieval results of state-of-the-art search engines. We used Google as our original search engine throughout the whole process. We developed MagicWand, a web platform that standardizes the training, execution, and evaluation pipeline for intelligent agents running on Android devices. It can spin up multiple Android emulators and connect with real devices on the server, where each device is controlled by the user on a web interface or ADB-equivalent command interfaces.
@@ 4.1. Test Data: ”How-to” WeWeb dataset
A new ”How-to” WeWeb dataset has been collected using MagicWand, which covers instruction extraction, on-device execution, and data annotation for reranking. This dataset is used to measure the end-to-end accuracy of our system.
The finalized dataset version contains queries and instructions for 17 different apps spanning 9 categories of the Google Play Store and an additional ”System” domain for the default system app. We selected a single ”How-to” search query in the System domain. For other domains, we recruit 6 students to propose about ten ”How-to” search queries per app. Using these queries, we retrieved valid web pages from the top 20 Google search results555For the ”Entertainment” domain, there are 18 search results that didn’t have valid web pages, but linked to YouTube videos; For the ”News” category, 1 search result wasn’t a valid web page..
@@
We hired Amazon Mechanical Turk workers to extract relevant instructions from candidate web pages for ”How-to” queries. Multiple workers analyzed each page to ensure data quality, with consensus needed on the extraction results to label as a positive case. The resulting data, detailed in Table 3, includes statistics on the number of search queries, result pages, and pages with valid instructions for each domain.
@@
With search queries and instructions extracted (firstly by humans and further verified by the extraction module) from pages returned by Google, tasks were set up on MagicWand to assess and record the execution of these instructions on Android devices. 16 workers manually read each instruction, executed the instructions on MagicWand, and labeled whether the ”How-to” queries could be completed based on the instructions on each page. Out of all pages evaluated, 168 pages are labeled as successful (i.e. positive) and 64 queries were associated with at least one positively labeled page. These human labels are adopted as our ground truth , which is assigned a value of 1 if a human annotator can follow the instructions on the page and complete the task, otherwise 0.
We captured the actions performed by the workers on the devices, along with their annotations, to the well formatted files. These documents included screenshots, JSON files that detailed the history of actions taken, and comprehensive screen recordings capturing the entire interaction sequence. Additionally, a specialized plugin integrated within the platform was employed to extract the UI context (see Section 3.2.1 having each action executed by the workers).
@@ 4.2. Evaluation Metrics
Common reranking metrics for evaluating model performance are used: mean reciprocal rank (MRR) of the top one retrieved documents(MRR@1), precision of the top one and five retrieved documents (P@1 and P@5) and normalized discounted cumulative gain (NDCG) of the top five retrieved documents (NDCG@5).
@@ 4.3.1. Extraction
The state-of-art LLM, GPT4 (OpenAI et al., 2023), has been used to extract instructions from a candidate web page. We set the temperature parameter as 0 to encourage direct extraction and minimize generative content. We also design a prompt containing a list of valid and relevant instruction descriptions. We also try cohere (Documentation, 2024) using the same prompt, which could output accurate citations from the input web page, and found there was no noticeable performance difference compared with GPT4.
@@ 4.3.2. Pre-training Reranking Models
Since the ”How-to” WeWeb dataset is small, we decided to use the whole dataset for testing; thus the reranking is a zero-shot learning problem. We created synthetic data to pre-train pairwise LTR models. The synthetic pre-training data - ”How-to” META-GUI dataset was created based on META-GUI (Sun et al., 2022), a multimodal dataset for task-oriented conversational agents running on Android. The original data set includes 1125 conversations, 4684 dialogue turns, and 18337 Android execution trajectories. It covers six apps: weather, calendar, search, taxi, restaurant, and hotel mobile apps. This dataset has been used to generate a new ”How-to” Meta-GUI dataset by the following process:
Given 62 few-shot samples(the paired instruction, execution, search query), GPT3.5 is prompted to summarize an execution trajectory with its dialog turn to generate a list of step-by-step instructions and a relevant ”How-to” search query. These serve as positive examples.
@@
This synthetic dataset significantly differs from real data and the testing set. However, we expect possessing some form of pre-training material, even if not ideal, is preferable to having none at all, particularly in the context of zero-shot learning scenarios. We split ”How-to” META-GUI dataset into training, validation, and test sets, then train the LTR models with the most suitable hyperparameters. Next, the trained models are applied to ”How-to” WeWeb dataset.
@@
Regarding the model details, TMLP is a fully connected layers of 96-dimensional input without normalization, followed by two transformer blocks with 384-dimensional features and a dropout rate of 0.1. A post-processing model takes the output of the transformer blocks and uses a linear layer to output the reranking score. For NeuralNDCG loss, its temperature is set as 1.0 and we use a ”powered relevancies” gain function , where x is the relevance score. For LambdaRank loss, we set its parameters = 10, and = 1.0. Both are optimized by the same training setting: an Adam optimizer with an initial learning rate of 0.001 is applied with a learning rate scheduler of size 50 and decreasing rate = 0.1. Both training epochs and early stopping patience parameters are set to 20 with NDCG@5 as the validation metric.
@@ 4.4. Methods compared
LR: This method uses a logistic regression based LTR model, pre-trained on a different synthetic dataset.
NeuralNDCG + TMLP: This method uses a transformer with multilayer perception for LTR, pre-trained using NeuralNDCG loss (Qin et al., 2020) on a different synthetic dataset.
LambdaLoss + TMLP: This method uses a transformer with multilayer perception for LTR, pre-trained using Lambda loss (Wang et al., 2018) on a different synthetic dataset.
@@ 5.1. Comparison with Baselines
The performance metrics of different models are shown in Table 5. The ”Oracle” model has the highest performance across all metrics, which is expected as it serves as the ideal benchmark. The baseline (i.e. Google) has the lowest performance, showing that all other models have improved the original ranking. The improvements over the baseline are statistically significant for LR, NeuralNDCG + TMLP, and LambdaLoss + TMLP models. Despite the simple zero-shot learning nature of the models used in our experiments, the results demonstrate adding verification and reranking into the retrieval process can significantly enhance the performance of a leading baseline search engine.
In comparison, -based rank performs worse than other reranking models. This shows reranking using a strong yet single feature is feasible but not good enough. Neural ranking models (TMLP) are better than LR probably because they can capture feature interrelationships. However, the difference is not very big, possibly due to the low quality of the synthetic data used for pre-training.

### v3

@@ 3.2.1. Runtime UI Context Detection
This module collects data that characterizes the device’s current User Interface (UI) context. The UI context consists of a UI screenshot and metadata of ”actionable regions”. An ”actionable region” refers to a region on the screen where actions, such as clicking or scrolling, can be performed. These regions are hierarchically organized and can be associated with several control properties, like textual description, coordinate, scrollable, etc. For example, a region could be a navigation bar, with a back button and title as embedded child regions. To simplify, we currently focus on leveraging the textual description of the control directly associated with an ”actionable region”, like title, text description, etc.
Sometimes, a region may not directly contain any textual description, but its child regions often possess meaningful textural information, which is crucial for accurate action prediction but is unfortunately missing. We designate a control distillation algorithm to fill in this missing information, depicted in figure 3. First, invisible or non-actionable regions are filtered out. Then, descriptions from the nearest non-actionable child regions are consolidated hierarchically to represent the missing information.
3.2.2. Multimodal Action Prediction Model
Following Yang et al. (2023b); Zhang and Zhang (2023); Sun et al. (2022), the Action Prediction Model prompts a multimodal LLM, GPT4-V OpenAI et al. (2023); Yang et al. (2023a) in our current setting, where the current UI context and multimodal action chains 333That is history of UI screenshots paired with each action taken are passed as input to choose the target action from a set of candidate actions 444Injecting the UI context into the prompt can help model better determine the next action based on our primary experiment. We expect more accurate action prediction results, if we can use this information appropriately as an input to train or finetune the model., such as click, swipe, input, back, etc.
For timestep t , GPT4-V will predict the next action given a prompt constructed using the following parameters:
@@ 3.2.2. Multimodal Action Prediction Model
: List of screenshots of UI screens from previous actions.
: Action description for each previous action, including the action name and the chosen actionable region description.
: Set of possible actions based on available properties of distilled controls in .
: Information not covered by instructions but required for agent execution. For instance, a specific username and password are needed, when the instruction is ”enter user name and password when open WeChat”.
@@ 3.2.3. Execution Proxy
It takes a predicted action and its parameters returned from the Action Prediction Model, locates the action to the corresponding region, and finally executes it. Considering a mobile application context can change dynamically, we introduce a heuristic to validate the predicted control associated with the target action and elegantly find a suitable alternative when mismatched. For flexibility, We also designate a fallback mechanism for candidate actions, so some of them can adjust slightly given different application settings. For instance, a click action can be tried with a similar gesture under instant Message application settings.
@@
It synchronizes the execution information of the mobile agents to external modules, helping to analyze and monitor the execution progress. In addition, this module also implements analysis plugins to integrate with GPT4V and GPT4 to summarize and evaluate the status of each execution trajectory.
@@ 3.3. Reranking
Execution information, both textual and visual information (UI screenshot and corresponding actionable regions), has been collected as listed in table 1 and converted into a reranking feature vector.
To avoid unnecessary complexity initially, we now use statistical measures of execution information, and leave it as future work to explore more advanced feature generation techniques, such as time series transformer or feature learning.
@@ 3.3.1. Feature Design
All textual and visual features are normalized to a float value between 0 and 1. Below is a brief description of them.
@@
Q is the set of tokens from the search query without stopword, where is the term frequency of token in the extracted instructions I.
@@
: Average term frequency of common keywords in instructions: This feature measures the average frequency of common keywords found in the extracted instructions. We have used instructions from about 1000 web pages from Google Android Help Website and identified a set of most common keywords, excluding stopwords.
@@
: ”Instructions completion degree” measures the degree to which the extracted instructions have been completed. GPT4-V is prompted to return a completion score of [0, 1], using the history of UI screenshots for each action taken during execution and extracted instructions as input.
: average, min, max, variance of Action term ratio in instructions measures the average, minimal, maximum, standard deviation of the percentage of instruction’s keyword that appears in each action description during the entire execution.
: average, min, max, variance of Visible UI Term Ratio in instructions measures the average, minimal, maximum, standard deviation of the percentage of the current screen UI control’s visible texts appeared in the instructions during the entire execution. This metric indicates the alignment between the current UI screen and the instructions. For example, for user registration, instructions often contain reference fields like username, password, and gender. A higher correspondence of these terms with UI elements suggests that the agent is more likely to move forward along the instructions.
: average, min, max, variance of Distilled UI ratio in instructions describes the average, minimal, maximum, standard deviation of the occurrence of distilled UI control text in instructions during the entire execution. These metrics reflect the likelihood of our execution model accurately selecting the appropriate UI control from all distilled options.
to : These capture how much of each extracted step is executed by comparing the list of actions taken for a given step555One step in a plan may require multiple actions to complete to its step text.
@@
Due to the page limit, we leave the exact equations for calculating the heuristic features to in our open-source website (TBA).
@@ 3.3.2. Reranking Models
In this section, we will illustrate which candidate pages have to be processed for reranking features and the details of reranking models.
@@
When the system fails to extract instructions or those pages do not contain any instruction, a candidate page does not advance to the on-device execution stage. After that, our reranking modules prioritize all verified pages using a heuristic. In our experiments, a page is considered as verified only if it progresses to stage 2 and satisfies the following criteria: the execution proxy must successfully execute at least one step and the predicted instruction completion rate is non-zero. All other Pages are kept in their original relative order (i.e., not reranked) and put after the verified pages.
@@
Without losing the generality, three different learning-to-rank(LTR) models have been trained to rerank the verified pages using the aforementioned features.
The first model is a simple pairwise LTR logistic regression model, which optimizes the relative order of each pair of pages. Another two types of models are multilayer perception (MLP) transformer models, short for TMLP, using two different loss functions. One is LambdaLoss (Wang et al., 2018), a pairwise loss function. For each pair of items, the lambda value represents the change in the overall ranking metric, if the order of the two items were swapped. The lambda values are used to weight the errors in the pairwise comparisons, thus this loss function will penalize misorderings more if they have a higher lambda value. Another is NeuralNDCG loss (Qin et al., 2020), a method directly optimizing NDCG metric. This method introduces a NeuralSort operator based on a differentiable function (softmax) to approximate the non-differentiable ground-truth sort operator.
@@ 4. Experiments
Experiments have been designed to assess whether the proposed method enhances the retrieval performance of state-of-the-art search engines. We used Google as our original search engine throughout the whole process. Our experiments are hosted on MagicWand, a web platform that standardizes the training, execution, and evaluation pipeline for intelligent agents running on Android devices. It can spin up multiple Android emulators or connect with real devices on the server, which can be controlled by the user on a web interface or ADB-equivalent command interface.
@@ 4.1. Test Data: ”How-to” WeWeb dataset
A new ”How-to” WeWeb dataset has been collected using MagicWand, which covers instruction extraction, on-device execution, and data annotation for reranking. This dataset is used to measure the end-to-end reranking accuracy.
Its finalized version contains queries and instructions for 17 different apps spanning 9 categories of the Google Play Store and an additional ”System” domain for the default system app 666Based on Google result, the default system app category has higher proportion of web pages with instructions in top 20 result. In comparison, other categories vary for instruction proportion and therefore deserve more attention. We selected a single ”How-to” search query in the System domain. For other domains, we recruit 6 students to propose about ten ”How-to” search queries per app. Using these queries, we retrieved valid web pages from the top 20 Google search results777For the ”Entertainment” domain, there are 18 search results that didn’t have any valid web page, but linked to YouTube videos; For the ”News” category, 1 search result wasn’t a valid web page..
@@
We hired Amazon Mechanical Turk workers to extract relevant instructions from candidate web pages for ”How-to” queries. Multiple workers analyzed each page to ensure data quality, requiring consensus on the extraction results to be labeled as positive. The resulting data, detailed in Table 3, includes statistics on the number of search queries, result pages, and pages with valid instructions for each domain.
@@
With search queries and instructions extracted (firstly by humans and further verified by the extraction module) from pages, tasks were set up on MagicWand to record and access the execution of instructions on Android devices. 16 workers read each instruction, executed the instructions manually, and labeled whether the assigned ”How-to” queries could be completed. Out of all pages evaluated, 168 pages are labeled as successful (i.e. positive) and 64 queries were associated with at least one positively labeled page. These human labels are used as the ground truth : it is assigned with a value of 1 if a human annotator can follow the instructions on the page and complete the task, otherwise 0.
We captured the actions performed by the workers, along with their annotations, and saved them into documents with a pre-defined format. These documents included screenshots, JSON files that detailed the history of actions taken, and screen recordings capturing the entire interaction sequence. Additionally, a specialized plugin integrated with the platform was employed to extract the UI context.
@@ 4.2. Evaluation Metrics
Common reranking metrics are used: mean reciprocal rank (MRR) of the top one retrieved documents(MRR@1), precision of the top one and five retrieved documents (P@1 and P@5) and normalized discounted cumulative gain (NDCG) of the top five retrieved documents (NDCG@5).
@@ 4.3.1. Extraction
The state-of-the-art LLM, GPT4 (OpenAI et al., 2023), is used to extract instructions from a candidate web page in our experiment, following Sarkhel et al. (2023); Sood and Dandapat (2023). We set the temperature parameter as 0 to encourage direct extraction with minimizing generative content. We also design a prompt containing a list of valid and relevant instruction descriptions. For comparison, we have also tested cohere (Documentation, 2024) using the same prompt, which could output accurate citations from the input web page, and found no noticeable performance difference compared with GPT4.
@@
4.3.2. On-Device Execution
Device setup
On MagicWand, we have reserved an Android emulator to run tasks given instructions and collect execution statistics 888Realistically, multiple emulators with various settings better simulate realistic cases. However, one device can help simplify our initial work and device maintenance.. For efficiency, a batch-supported execution engine based on adbutils (adb library for adb service, 2022) is also prepared to run multiple tasks on different devices simultaneously.
Multimodal action prediction model
To avoid invalid action parameters and keep exploring capacity, GPT4-V parameter ”temperature” is set to 0.3 together with ”max_tokens” of 300. We specify instructions, historical actions and UI screenshots in the prompt and adjust the prompt according to distilled UI controls.
Execution proxy
The UI context representation data with the execution status are saved into the central storage with a JSON index file describing the execution and associated resources. A GPT4V python plugin was developed to evaluate the execution information and generate the task completion rate for each execution.
@@ 4.3.3. Pre-training Reranking Models
As the ”How-to” WeWeb dataset is small, we decided to use it for testing; thus, the reranking is a zero-shot learning problem. We created another synthetic data to pre-train pairwise LTR models. The synthetic pre-training data - ”How-to” META-GUI dataset was created based on META-GUI (Sun et al., 2022), a multimodal dataset for task-oriented conversational agents running on Android. The original data set includes 1125 conversations, 4684 dialogue turns, and 18337 Android execution trajectories. It covers six apps: weather, calendar, search, taxi, restaurant, and hotel mobile apps. This dataset has been used to generate a new ”How-to” Meta-GUI dataset by the following process:
Given 62 few-shot samples(the paired instruction, execution, search query), GPT3.5 is prompted to summarize an execution trajectory with its dialog turn to generate a list of step-by-step instructions and a relevant ”How-to” search query. These are positive examples.
@@
This synthetic dataset significantly differs from real data and the testing set. However, we expect possessing some form of pre-training material, even if not ideal, is preferable to having none at all, particularly in the zero-shot learning settings. We split ”How-to” META-GUI dataset into training, validation, and test sets, then train the LTR models with the most suitable hyperparameters. Next, the trained models are applied to ”How-to” WeWeb dataset.
@@
Regarding the model details, TMLP is a fully connected layer of 96-dimensional input without normalization, followed by two transformer blocks with 384-dimensional features and a dropout rate of 0.1. A post-processing model takes the output of the transformer blocks and uses a linear layer to output the reranking score. For NeuralNDCG loss, its temperature is set as 1.0 and we use a ”powered relevancies” gain function , where x is the relevance score. For LambdaRank loss, we set its parameters = 10, and = 1.0. They are all optimized in the same training setting: an Adam optimizer with an initial learning rate of 0.001 is applied with a learning rate scheduler of size 50 and a decreasing rate = 0.1. Both training epochs and early stopping patience parameters are set to 20 with NDCG@5 as the validation metric.
@@ 4.4. Methods compared
LR: This method uses a logistic regression-based LTR model, pre-trained on the synthetic dataset.
NeuralNDCG + TMLP: This method uses a transformer with multilayer perception for LTR, pre-trained using NeuralNDCG loss (Qin et al., 2020) on the synthetic dataset.
LambdaLoss + TMLP: This method uses a transformer with multilayer perception for LTR, pre-trained using Lambda loss (Wang et al., 2018) on the synthetic dataset.
@@ 5.1. Comparison with Baselines
The performance metrics of different models are shown in Table 5. The ”Oracle” model has the highest performance across all metrics, which is expected as the ideal benchmark. The baseline (i.e. Google) has the lowest performance, showing that all other models have improved the original Google ranking when taking execution status into consideration. The improvements over the baseline are statistically significant on LR, NeuralNDCG + TMLP, and LambdaLoss + TMLP models. Despite their zero-shot learning nature, the results demonstrate adding verification and reranking into the retrieval process can significantly enhance the performance of a leading baseline search engine.
In comparison, -based rank performs worse than other reranking models. This shows reranking using a strong yet single feature is feasible but not good enough. Two neural ranking models (TMLP) are better than LR probably because they can capture feature interrelationships. However, the difference is not very big, possibly due to the low quality of the synthetic data used for pre-training.